How to solve
48 divided 4x5 - 2x5 =

Answer by DRichardson(63)

Hello There,
48 divided 4x5 - 2x5 =
this is a pretty simple one
i believe it should be written like this
48 / (4 x 5) - ( 2 x 5)=???
48/ ( 20) - (10)
2.4 - 10 = -7.6

If not then i will solve it also how you wrote it:
290 would be your answer.
If you need to and are allowed, try using a calculator for help on tricky equations :)
Delilah R